OpenStudy (anonymous):

What is the first step in solving and graphing the inequality 2x-6y>8 ? A. Subtract 2x from both sides of the inequality. B. Test the point (0, 0). C. Switch the inequality to "less than or equal to." D. Graph the inequality.

OpenStudy (mathmale):

Typically one would isolate the ' y ' term on one side of the inequality, but there's no law that says you can't do the alternative: isolate ' x ' on one side. If 2x - 6y > 8 try isolating y. This sort of problem is easier if you start out by obtaining a POSITIVE coefficient for y, so I'd suggest you add 6y to both sides of this inequality. What do you get, and what's the next step?

OpenStudy (anonymous):

A ? @mathmale

OpenStudy (mathmale):

I'm really anxious to help. Help me to help you by sharing the work you've done towards choosing the correct answer, OK?

OpenStudy (anonymous):

unless i can plug in the ( 0 , 0 ) . && Ok .

OpenStudy (mathmale):

If 2x - 6y > 8 try isolating y. This sort of problem is easier if you start out by obtaining a POSITIVE coefficient for y, so I'd suggest you add 6y to both sides of this inequality. What do you get, and what's the next step? 2x - 6y > 8 + 6y = +6y ---------- ----- complete this work, please.

OpenStudy (anonymous):

what to do after 2x>14y or did i do this wrong ?

OpenStudy (mathmale):

Your 2x on the left is fine. But combining 8 with 6y does not result in 14y. Mind trying that again?

OpenStudy (mathmale):

The ' 8 ' and the ' 6y ' are NOT "like" terms, and thus cannot be combined. 8y+6y=14y (yes) 8 +6y = 14y NO So, please re-write the inequality: 2x > ???

OpenStudy (anonymous):

so it would be , 2x > 6y + 8 .